# ADAPT: A Pseudo-labeling Approach to Combat Concept Drift in Malware Detection

### v1

@@ VI-D Evaluation
VII Results and Discussions
TABLE II: Performance on the Drebin dataset.
Method F1 FPR FNR
Offline ML RF 33.2±1.00 0.14±0.04 78.7±0.71
XGBoost 44.9±0.00 0.66±0.00 65.6±0.00
MLP 40.9±1.10 0.57±0.04 69.8±0.82
Self-Training ARF 34.0±7.14 9.88±7.24 45.8±5.21
DE++ 34.7±1.89 7.84±1.17 47.1±3.08
Insomnia 49.8±0.03 0.70±0.01 60.6±0.15
MORSE 41.2±0.66 4.50±0.15 54.7±0.96
Proposed RF+ADAPT 49.8±0.26 0.21±0.01 64.3±0.23
XGBoost+ADAPT 57.1±0.37 0.24±0.04 56.6±0.37
MLP+ADAPT 50.5±2.40 0.77±0.49 59.1±0.96
Figure 7: Performance on the Drebin dataset: F1 score (left), FPR (middle), and FNR (right) over test months.
VII-A Drebin Dataset
Table II presents the results on the Drebin dataset for various methods. Drebin is the most challenging dataset, as reflected by the poor F1 scores of the offline machine-learning models. Among the four baselines, ARF, DE++, and MORSE fail to show notable improvement over the offline models. In contrast, Insomnia performs better than the offline MLP and XGBoost models, indicating that our adaptation of Insomnia for the training process is effective for Android malware detection under concept drift, even though Insomnia was originally designed for network intrusion detection.
Our proposed semi-supervised method shows considerable improvements across all baseline models. The absolute improvement in F1-score for RF, XGBoost, and MLP is 16.6%, 12.2%, and 9.6%, respectively. This improvement is primarily attributed to a significant reduction in false negatives (i.e., malware samples misclassified as benign). Notably, our method does not increase the false positive rate; in fact, it reduces the false positive rate by 63.6% compared to the baseline XGBoost model.
Figure 7 shows the F1 score over the test months for four different models on the Drebin dataset. Our proposed method consistently outperforms the other models across most test months. As evident from the figure, all three adaptation algorithms can mitigate the effects of concept drift to varying degrees.
TABLE III: Performance on the APIGraph dataset.
Method F1 FPR FNR
Offline ML RF 46.8±3.89 0.34±0.08 66.9±3.76
XGBoost 66.5±0.90 0.97±0.00 42.7±0.90
MLP 57.0±3.52 1.67±0.64 49.9±6.78
Self-Training ARF 61.1±5.84 5.28±3.68 25.0±10.8
DE++ 42.3±30.1 0.33±0.25 65.6±24.6
Insomnia 68.1±0.43 2.05±0.10 31.8±1.00
MORSE 75.3±0.82 1.82±0.23 22.2±0.77
Proposed RF+ADAPT 67.2±0.29 0.34±0.01 46.0±0.36
XGBoost+ADAPT 73.2±0.92 2.26±0.23 22.2±0.48
MLP+ADAPT 76.8±0.52 0.81±0.10 29.4±0.97
@@ VII-A Drebin Dataset
VII-B APIGraph Dataset
The results for the APIGraph dataset are presented in Table III. The baseline machine learning models perform better on this dataset, as its features are designed to be more resistant to concept drift [77]. Among the baselines, three self-training models, excluding DE++, show improvements over the offline models.
Our proposed method demonstrates significant improvements over the baseline models, particularly for RF and MLP, both of which exhibit around 20% increase in F1-score. While MORSE achieves the best False Negative Rate, our method achieves a lower False Positive Rate with the best-performing MLP model.
Figure 8 shows the F1-score over the test months for four models on the APIGraph dataset. The performance of all three adaptation models remains relatively stable for the first two years of test data, after which Insomnia shows a sharp decline. Our proposed method and MORSE perform similarly for most of the test period, except towards the end, where more significant drift occurs, and our proposed method outperforms MORSE under these conditions.
TABLE IV: Performance on BODMAS dataset.
Method F1 FPR FNR
Offline ML RF 96.3±0.30 0.21±0.01 6.69±0.55
XGBoost 99.2±0.00 0.30±0.00 1.23±0.00
MLP 98.2±0.32 0.64±0.15 2.89±0.69
Self-Training ARF 97.5±0.00 2.13±0.00 2.28±0.00
DE++ 90.78±1.0 15.1±1.88 1.83±0.69
Insomnia 97.2±0.28 0.56±0.05 4.81±0.51
MORSE 98.8±0.12 0.48±0.06 1.73±0.32
Proposed RF+ADAPT 96.1±0.42 0.22±0.01 7.11±0.75
XGBoost+ADAPT 99.6±0.02 0.24±0.01 0.61±0.03
MLP+ADAPT 98.6±0.25 0.41±0.02 2.26±0.50
TABLE V: Performance on EMBER dataset.
Method F1 FPR FNR
Offline ML RF 86.3±0.29 3.47±0.02 21.7±0.45
XGBoost 92.3±0.00 3.58±0.00 11.4±0.00
MLP 88.1±0.28 5.66±0.50 17.1±0.39
Self-Training ARF 69.7±0.00 91.8±0.00 1.78±0.00
DE++ 76.9±0.82 24.9±0.73 22.8±1.21
Insomnia 87.1±0.35 6.06±0.13 18.5±0.66
MORSE 89.8±0.22 6.34±0.16 13.7±0.36
Proposed RF+ADAPT 86.2±0.26 3.56±0.03 21.7±0.42
XGBoost+ADAPT 93.5±0.05 4.45±0.07 8.70±0.11
MLP+ADAPT 89.1±0.37 5.01±0.18 16.0±0.73
VII-C Windows Datasets
@@
The results for the BODMAS and EMBER datasets are presented in Table IV and Table V, respectively. The baseline models achieve significantly higher F1 scores on these datasets, indicating a less pronounced impact of concept drift. The performance on the EMBER dataset is slightly lower, primarily because we did not perform separate hyperparameter tuning for this dataset and used only one month of data for training, compared to the three months used for BODMAS.
None of the self-training baseline methods on the BODMAS dataset outperform the baseline XGBoost model, which already achieves an F1-score of 99.2%. However, our proposed method improves this score by an additional 0.4%, with reductions in both the False Positive Rate (FPR) and False Negative Rate (FNR). This underscores the importance of our model-agnostic approach, as tree-based models may inherently perform better than neural networks on certain datasets [33]. A similar improvement is observed for MLP, although the performance of RF decreases—a trend that is also consistent with the results on the EMBER dataset.
Since the BODMAS dataset does not exhibit significant drift, we curated an additional variant by limiting the number of malware families in the training data to evaluate model performance under more severe drift. This setting, described in Appendix F, demonstrates that ADAPT exhibits greater resilience to distributional shifts.
@@
For the EMBER dataset, the baseline XGBoost model outperforms all self-training methods. Our method further improves upon the XGBoost baseline by 1.2%. Among the baseline self-training models, DE++ and ARF perform substantially worse than the baseline models, suggesting they are more susceptible to self-poisoning and highly sensitive to threshold settings on these datasets.
TABLE VI: Performance on PDF malware dataset.
Method F1 FPR FNR
Offline ML RF 98.0 ± 0.87 0.07 ± 0.00 3.31 ± 1.57
XGBoost 97.9 ± 0.95 0.14 ± 0.04 3.17 ± 1.50
MLP 97.4 ± 0.17 0.11 ± 0.01 4.26 ± 0.30
Self-Training ARF 29.7 ± 0.00 0.01 ± 0.00 80.0 ± 0.00
DE++ 60.8 ± 10.5 19.3 ± 6.70 0.28 ± 0.06
Insomnia 88.2 ± 7.41 0.09 ± 0.02 15.4 ± 8.87
MORSE 97.4 ± 0.13 0.12 ± 0.01 4.16 ± 0.22
Proposed RF+ADAPT 99.1 ± 0.11 0.09 ± 0.00 1.26 ± 0.20
XGBoost+ADAPT 98.6 ± 0.36 0.10 ± 0.00 2.03 ± 0.67
MLP+ADAPT 97.2 ± 0.12 0.09 ± 0.01 4.73 ± 0.24
VII-D PDF Dataset
Table VI presents the results on the PDF malware dataset. Interestingly, Random Forest outperforms both XGBoost and MLP on this dataset, suggesting that more complex models may be prone to overfitting. Nonetheless, our proposed method, ADAPT, improves the F1-score by 1.1%, primarily due to a reduction in the False Negative Rate (FNR). This result again highlights the effectiveness of our model-agnostic approach, which can enhance performance across different classifiers.
TABLE VII: Difference between False Positive Rate (FPR) and False Negative Rate (FNR) across different models and datasets. Positive values indicate a reduction in error, while negative values indicate an increase in error using the proposed ADAPT algorithm.
@@ VII-D PDF Dataset
VII-E Effectiveness of ADAPT Across Models
Although MLP+ADAPT achieves the highest performance over the full six-year test period on the APIGraph dataset, XGBoost+ADAPT (83.7% F1) outperforms both MLP+ADAPT (82.7% F1) and MORSE (80.5% F1) in the first year of test data, a timeframe that more realistically represents practical deployment scenarios for unlabeled adaptation. This highlights that while ADAPT is model-agnostic, its strong synergy with XGBoost enhances its versatility as a broadly applicable solution. A statistical analysis of ADAPT, provided in Appendix E, further substantiates this finding.
VII-F Error Analysis
Table VII summarizes the changes in False Positive Rate (FPR) and False Negative Rate (FNR) between baseline models and those trained with ADAPT. Across datasets, F1-score improvements are mainly due to reduced FNR, except for Random Forest on BODMAS and MLP on PDF. This suggests that ADAPT enables the model to better learn from uncertain malware samples, improving its ability to correctly classify similar cases upon retraining. In most scenarios, FPR is maintained or reduced alongside FNR, though exceptions exist—e.g., XGBoost on APIGraph, which shows decreased FNR but increased FPR. Neural networks generally outperform other models in reducing FPR. We further examine this in the context of pseudo-labeling errors on APIGraph.
TABLE VIII: Performance metrics (F1, FPR, FNR) for RF, XGBoost, and MLP with various components removed on the first six months of the APIGraph test set.
Component Removed RF XGBoost MLP
@@ VII-F Error Analysis
VII-G Ablation Study
We conduct an ablation study by removing different components of ADAPT on the APIGraph dataset, as shown in Table VIII for the first six months of test data. In this experiment, we replace the class-specific adaptive thresholding scheme with a fixed threshold of $\tau_{b}$. We observe that each component contributes to the overall improvement in F1-score and FNR, albeit varying degrees across different models. While replacing adaptive thresholding reduces the FPR marginally due to the higher threshold, as discussed in Section VII-F, it leads to a significant increase in the FNR, as drifted malware samples are not included in the retraining phase.
TABLE IX: Performance on Drebin and APIGraph with active learning (AL) with 50 monthly annotation budget.
@@ VIII Compatibility with Active Learning
Table IX presents the results with a monthly annotation budget of 50 on the Android datasets. For the XGBoost baseline model, we randomly select the same number of annotated samples per month to ensure consistency with the active learning methods following [11]. We report active learning results based on uncertainty sampling, using XGBoost and the Hierarchical Contrastive Learning Classifier (HCC) introduced in [23]. HCC, the state-of-the-art method for active learning with neural networks, leverages contrastive learning with a hierarchical loss function to distinguish between samples from different malware families in the embedding space.
From Table IX, we observe that active learning improves performance over random selection for the XGBoost model. Furthermore, ADAPT enhances this performance by leveraging the remaining unlabeled data. Specifically, we observe a 0.4% increase in F1-score for the Drebin dataset and a 0.7% improvement for APIGraph. While these gains are relatively modest, they indicate that ADAPT can effectively benefit from active learning. Although our current approach involves a straightforward integration with active learning, exploring more sophisticated applications remains an avenue for future work.
IX Extension to Multiclass Classification
Datasets. We follow the same temporal split as in the binary classification task for training, validation, and test splits in both datasets. We select the top 10 most frequent malware families for each dataset in the training split and include the benign class, resulting in an 11-class classification task. We perform a hyperparameter search on the BODMAS dataset using 100 iterations per model and apply the selected hyperparameters directly to the EMBER dataset without further tuning. Appendix C provides additional details on these datasets.
Algorithm Adjustments. In the binary detection setting, we use separate threshold parameters, $\tau_{b}$ for benign samples and $\tau_{m}$ for malware samples, updated dynamically based on model predictions on unlabeled data using the adaptation parameter $\lambda$. While $\lambda$ is shared across all classes, introducing separate thresholds for each malware family in the multiclass setting would significantly increase the number of parameters, complicating hyperparameter tuning. To mitigate this, we share $\tau_{m}$ across all malware classes, keeping the number of hyperparameters the same as in the binary setting (five), regardless of the number of malware families. Appendix C explores a scenario without a benign class, where only malware families are classified. Apart from this threshold adjustment, the rest of the algorithm, including augmentation and mixup, remains unchanged.
TABLE X: Performance on multiclass classification (11-class) on BODMAS and EMBER datasets.
@@ IX Extension to Multiclass Classification
Results. We report the macro F1-score, precision, and recall in Table X for the XGBoost baseline, other self-training methods, and XGBoost+ADAPT on the two datasets. While ADAPT improves over the baseline on both datasets, other self-training methods perform worse than the XGBoost baseline. This effect is particularly pronounced in the EMBER dataset, which is more challenging and where we do not perform additional hyperparameter tuning. Notably, ADAPT improves the average F1-score over XGBoost by 2.8%, whereas other self-training methods perform significantly worse than the baseline, suggesting that ADAPT is more robust to hyperparameter values compared to other methods on EMBER feature.
TABLE XI: Model performance as unknown malware from new families is introduced during testing.
Metric Additional new families
5 10 20
Evasion success rate (%) 9.83 7.10 6.59
AUC 0.706 0.713 0.713
Unknown Family Detection. To evaluate the open-set performance of the multi-class malware classifier, we conducted a monthly analysis with varying levels of novelty in the unknown samples. For each month, we selected the top-$k$ most frequent unseen malware families ($k\in\{5,10,20\}$) and computed two key metrics: the evasion success rate, defined as the fraction of unknown samples misclassified as benign, and the AUC, which reflects the model’s ability to separate known from unknown samples based on confidence scores. This AUC metric aligns with out-of-distribution (OOD) detection [71], where, ideally, lower confidence is assigned to unseen classes. Table XI presents the average results on the EMBER dataset using the XGBoost+ADAPT model, highlighting its tendency to misclassify unknown malware as known malicious classes.
TABLE XII: Performance on multiclass classification on BODMAS and EMBER datasets with active learning.
@@
Active Learning. We conduct additional experiments using an annotation budget of 50 samples per month for active learning with the XGBoost model. Table XII presents the results for random sample selection and active learning with the 50 most uncertain samples. Additionally, applying ADAPT with active learning improves performance over the random baseline by 3% on the BODMAS dataset and 2.6% on the EMBER dataset.

### v2

@@ VI-D Evaluation
VII Results and Analysis
@@
Figure 6: Performance on the Drebin dataset: F1 score (left), FNR (middle), and FPR (right) over test months.
Figure 7: Performance on the BODMAS dataset: F1 score (left), FNR (middle), and FPR (right) over test months.
Figure 8: Performance on the EMBER dataset: F1 score (left), FNR (middle), and FPR (right) over test months.
Figure 9: Performance on the PDF dataset: F1 score (left), FNR (middle), and FPR (right) over test months.
Table II presents the results of various methods across five datasets. XGBoost+ADAPT achieves the highest overall F1-score on the Drebin, BODMAS, and EMBER datasets, while MLP+ADAPT attains the best performance on the APIGraph dataset. Random Forest+ADAPT yields the top result on the PDF malware dataset. Across all benchmarks, ADAPT consistently enhances the performance of baseline machine learning models, with particularly notable gains on the Android malware dataset, where concept drift causes greater performance degradation.
We also report results for an Oracle setting, where ground-truth labels are available for retraining, but augmentation and mixup components used in ADAPT are omitted. The results indicate that, for the Drebin and APIGraph datasets, the F1-score of ADAPT lags behind the Oracle model by more than 10%. However, this gap is less pronounced on the other three datasets, and on the EMBER dataset, XGBoost + ADAPT even surpasses the Oracle, likely due to the benefits introduced by data augmentation and mixup.
Drebin Dataset
Drebin is the most challenging dataset, as evidenced by the consistently low F1-scores of the offline machine learning models. Among the four self-training baselines, ARF, DE++, and MORSE do not yield notable improvements over the offline models in terms of F1-score. In contrast, Insomnia outperforms the offline MLP and XGBoost models, highlighting the effectiveness of our Insomnia adaptation for malware detection under concept drift. ADAPT achieves substantial improvements across all baseline models, with absolute F1-score gains of 16.6%, 12.2%, and 9.6% for RF, XGBoost, and MLP, respectively. This improvement is primarily attributed to a significant reduction in false negatives. Notably, our method does not increase the false positive rate; in fact, it reduces the false positive rate by 63.6% compared to the baseline XGBoost model.
Figure 6 illustrates the temporal dynamics of F1, FPR, and FNR over the test months for four models on the Drebin dataset. Notably, the XGBoost+ADAPT method consistently outperforms the other models in F1-score across most test months and is particularly effective at mitigating the sharp performance drop experienced by the baseline XGBoost model in certain months, such as month 9 (2021-03).
APIGraph Dataset
The baseline machine learning models achieve higher performance on the APIGraph dataset compared to Drebin, as its features are specifically designed to be more resilient to concept drift [83]. Among the self-training baselines, all except DE++ show improvements over the offline models. Our proposed ADAPT method further improves upon the baselines, particularly for RF and MLP, both of which exhibit an approximately 20% increase in F1-score. While MORSE achieves the lowest False Negative Rate, the MLP+ADAPT method attains a lower False Positive Rate along with the best overall F1-score.
Figure 5 illustrates the F1-score over the test months for four models on the APIGraph dataset. The performance of all three adaptation models remains relatively stable during the first two years of test data, after which Insomnia experiences a sharp decline. ADAPT and MORSE perform similarly throughout most of the test period, except towards the end where more significant drift occurs.
@@
The baseline models achieve notably high F1-scores on the BODMAS dataset, indicating that concept drift has a less pronounced effect in this setting. None of the self-training baseline methods surpasses the performance of the baseline XGBoost model, which already attains an F1-score of 99.2%. Our proposed ADAPT method further increases this score by 0.4%, while also reducing both the False Positive Rate (FPR) and False Negative Rate (FNR). This result underscores the value of our model-agnostic approach, as tree-based models such as XGBoost may inherently outperform neural networks on certain datasets [36]. A similar improvement is observed for MLP, although the performance of RF decreases. Figure 7 presents the monthly results for different models, highlighting that XGBoost+ADAPT consistently maintains high performance throughout the test period.
Figure 10: Exposure over test months on different datasets.
@@
Performance on the EMBER dataset is slightly lower, mainly due to the use of only one month of training data and the absence of separate hyperparameter tuning, in contrast to the three months of training used for BODMAS. On EMBER, the baseline XGBoost model outperforms all self-training methods, but our proposed ADAPT method still achieves a further 1.2% improvement in F1-score over the XGBoost baseline. Among the self-training baselines, DE++ and ARF perform significantly worse than the offline models, indicating greater susceptibility to self-poisoning and a high sensitivity to threshold settings in this dataset. Figure 8 shows the monthly performance of different models, highlighting that XGBoost+ADAPT consistently achieves the highest performance throughout the test period.
PDF Dataset
Figure 9 shows the weekly performance of different models on the PDF malware dataset. Random Forest outperforms both XGBoost and MLP, suggesting that more complex models may be susceptible to overfitting in this setting. Nonetheless, our proposed method, ADAPT, improves the F1-score by 1.1%, primarily through a reduction in the False Negative Rate (FNR).
VII-A Exposure Analysis
To better capture the real-world risk associated with missed detections, we evaluate models using the Absolute Exposure (AE) metric introduced in [19]. AE quantifies the cumulative number of false negatives over time, reflecting the duration and extent to which users remain exposed to undetected threats. Unlike traditional accuracy metrics, AE directly measures how long malicious samples evade detection, providing a more practical assessment of security risk in streaming or temporally evolving scenarios.
Figure 10 presents the Absolute Exposure values for baseline and ADAPT models across all datasets. Notably, XGBoost+ADAPT achieves the lowest AE in every dataset except for the PDF malware dataset, where Random Forest + ADAPT attains the best performance.
VII-B Effectiveness of ADAPT Across Models
Although MLP+ADAPT achieves the highest performance over the full six-year test period on the APIGraph dataset, XGBoost+ADAPT (83.7% F1) outperforms both MLP+ADAPT (82.7% F1) and MORSE (80.5% F1) in the first year of test data, a timeframe that more realistically represents practical deployment scenarios for unlabeled adaptation. This highlights that while ADAPT is model-agnostic, its strong synergy with XGBoost enhances its versatility as a broadly applicable solution. The statistical analysis of ADAPT, provided in Section VII-E further substantiates this finding.
VII-C Error Analysis
Table III summarizes the changes in False Positive Rate (FPR) and False Negative Rate (FNR) between baseline models and those trained with ADAPT. Across datasets, F1-score improvements are mainly due to reduced FNR, except for Random Forest on BODMAS and MLP on PDF. This suggests that ADAPT enables the model to better learn from uncertain malware samples, improving its ability to correctly classify similar cases upon retraining. In most scenarios, FPR is maintained or reduced alongside FNR, though exceptions exist—e.g., XGBoost on APIGraph, which shows decreased FNR but increased FPR. Neural networks generally outperform other models in reducing FPR. We further examine this in the context of pseudo-labeling errors on APIGraph.
TABLE III: Difference between False Positive Rate (FPR) and False Negative Rate (FNR) across different models and datasets. Positive values indicate a reduction in error, while negative values indicate an increase in error using the ADAPT.
@@ VII-C Error Analysis
TABLE IV: Performance metrics (F1, FPR, FNR) for RF, XGBoost, and MLP with various components removed on the first six months of the APIGraph test set.
Component
Removed
RF XGBoost MLP
@@
VII-D Ablation and Mechanism Analysis
We conduct an ablation study by removing different components of ADAPT on the APIGraph dataset, as shown in Table IV for the first six months of test data. In this experiment, we replace the class-specific adaptive thresholding scheme with a fixed threshold of $\tau_{b}$. We observe that each component contributes to the overall improvement in F1-score and FNR, albeit varying degrees across different models. While replacing adaptive thresholding reduces the FPR marginally due to the higher threshold, as discussed in Section VII-C, it leads to a significant increase in the FNR, as drifted malware samples are not included in the retraining phase.
TABLE V: Feature Correlations Before and After Augmentation
Dataset Mean $|\rho(X)|$ Spearman ($\rho(X)$, $\rho(X_{\text{aug}})$)
APIGraph 0.0238 0.9247
Drebin 0.0269 0.9911
Bodmas 0.0276 0.9491
PDF 0.0857 0.9975
Data Augmentation & Consistency. Our feature-space augmentation does not guarantee valid samples, as modeling complex feature dependencies is intractable. To assess the impact, we analyze whether augmentation disrupts feature relationships by computing the mean absolute correlation in the original data and the Spearman correlation between original and augmented correlation matrices. As shown in Table V, low mean correlations indicate weak dependencies, and high Spearman values confirm that feature relationships are largely preserved, suggesting that the overall statistical structure remains stable.
Effect of Mixup. Figure 12 shows the effect of Mixup on calibration for XGBoost on APIGraph. The scatter plots compare accuracy versus confidence for three models: base (no augmentation or Mixup), with augmentation only, and with both augmentation and Mixup. The base model is notably overconfident, while data augmentation improves calibration, and combining it with Mixup provides the closest alignment between confidence and accuracy, especially for high-confidence predictions.
VII-E Statistical Significance
Table VI presents the Wilcoxon signed-rank test p-values for comparing different machine learning models (Random Forest, MLP, and XGBoost) across multiple malware datasets. The Wilcoxon test [75] is a non-parametric statistical test used to assess whether two paired samples come from the same distribution. In this case, it evaluates whether the adapted model’s F1-scores significantly differ from the baseline in the test months. Cells are shaded in gray if the p-value is less than 0.05, indicating statistically significant differences. Notably, we observe consistently low p-values for the XGBoost models across all datasets, suggesting a significant effect of adaptation on their performance.
TABLE VI: Wilcoxon $p$-values for different models across datasets. Cells with $p<0.05$ are shaded.
Dataset Random Forest MLP XGBoost
Drebin $7.63\times 10^{-6}$ $5.34\times 10^{-5}$ $1.53\times 10^{-5}$
APIGraph $1.64\times 10^{-12}$ $1.64\times 10^{-12}$ $1.23\times 10^{-10}$
BODMAS 0.68750.68750.6875 0.031250.031250.03125 0.031250.031250.03125
EMBER 0.76460.76460.7646 0.032230.032230.03223 0.001950.001950.00195
PDF 0.43750.43750.4375 0.156250.156250.15625 0.031250.031250.03125
TABLE VII: Performance on Drebin and APIGraph with active learning (AL) with 50 monthly annotation budget.
@@ VIII Compatibility with Active Learning
Table VII presents the results with a monthly annotation budget of 50 on the Android datasets. For the XGBoost baseline model, we randomly select the same number of annotated samples per month to ensure consistency with the active learning methods following [10]. We report active learning results based on uncertainty sampling, using XGBoost and the Hierarchical Contrastive Learning Classifier (HCC) introduced in [24]. HCC, the state-of-the-art method for active learning with neural networks, leverages contrastive learning with a hierarchical loss function to distinguish between samples from different malware families in the embedding space.
From Table VII, we observe that active learning improves performance over random selection for the XGBoost model. Furthermore, ADAPT enhances this performance by leveraging the remaining unlabeled data. Specifically, we observe a 0.4% increase in F1-score for the Drebin dataset and a 0.7% improvement for APIGraph. While these gains are relatively modest, they indicate that ADAPT can effectively benefit from active learning. Although our current approach involves a straightforward integration with active learning, exploring more sophisticated applications remains an avenue for future work.
TABLE VIII: Performance on multiclass classification (11-class) on BODMAS and EMBER datasets.
@@
TABLE IX: Performance on multiclass classification on BODMAS and EMBER datasets with active learning.
@@
IX Extension to Multiclass Classification
Datasets. We follow the same temporal split as in the binary classification task for training, validation, and test splits in both datasets. We select the top 10 most frequent malware families for each dataset in the training split and include the benign class, resulting in an 11-class classification task. We perform a hyperparameter search on the BODMAS dataset using 100 iterations per model and apply the selected hyperparameters directly to the EMBER dataset without further tuning. Appendix B provides additional details on these datasets.
Algorithm Adjustments. In the binary detection setting, we use separate threshold parameters, $\tau_{b}$ for benign samples and $\tau_{m}$ for malware samples, updated dynamically based on model predictions on unlabeled data using the adaptation parameter $\lambda$. While $\lambda$ is shared across all classes, introducing separate thresholds for each malware family in the multiclass setting would significantly increase the number of parameters, complicating hyperparameter tuning. To mitigate this, we share $\tau_{m}$ across all malware classes, keeping the number of hyperparameters the same as in the binary setting (five), regardless of the number of malware families. Appendix B explores a scenario without a benign class, where only malware families are classified. Apart from this threshold adjustment, the rest of the algorithm, including augmentation and mixup, remains unchanged.
Results. We report the macro F1-score, precision, and recall in Table VIII for the XGBoost baseline, other self-training methods, and XGBoost+ADAPT on the two datasets. While ADAPT improves over the baseline on both datasets, other self-training methods perform worse than the XGBoost baseline. This effect is particularly pronounced in the EMBER dataset, which is more challenging and where we do not perform additional hyperparameter tuning. Notably, ADAPT improves the average F1-score over XGBoost by 2.8%, whereas other self-training methods perform significantly worse than the baseline, suggesting that ADAPT is more robust to hyperparameter values compared to other methods on EMBER feature.
[…]
Active Learning. We conduct additional experiments using an annotation budget of 50 samples per month for active learning with the XGBoost model. Table IX compares results for random sampling versus selecting the 50 most uncertain samples each month. Notably, integrating ADAPT with active learning improves performance over the random baseline by 3% on the BODMAS dataset and 2.6% on the EMBER dataset.